OpenStudy (mateeldaa):

I have this question for Discrete Math. We are currently on Simple Interest, but I have no idea how to solve this.: Given that u and v are real numbers and that 3^u=10 and 3^v=(1/2) check all the answers that are true statements. (3^u means 3 raised to the u power). 3^(uv)=5 3^(v+2)=18 3^(u+v)=5 3^(2-v)=18

OpenStudy (anonymous):

\[3^{v+2}=3^v\times 3^2=\frac{1}{2}\times 9\] so not that one

OpenStudy (anonymous):

\[3^{u+v}=3^u\times 3^v=10\times \frac{1}{2}=5\] so that looks good

OpenStudy (anonymous):

\[3^{2-v}=3^2\times 3^{-v}=9\times \frac{1}{3^v}=9\times 2=18\]

OpenStudy (anonymous):

Here are some identities you need to solve your question: \[a^{vu} = (a^{v})^{u}\]\[a^{u+v} = a^{u} \times a^{v}\]

OpenStudy (mateeldaa):

Thanks!! I'm not sure how to give you both medals for your help. :\

OpenStudy (anonymous):

don't worry about it lol.